OpenStudy (anonymous):

There were 48 runners to start a race. In the first half of the race, 1/3 of them dropped out. In the second half of the race, 3/4 of the remaining runners dropped out. How many runners finished the race?

OpenStudy (anonymous):

what is one third of 48?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

ok and what is 48-16?

OpenStudy (anonymous):

32 but i thought you had to make the fractions equal?

OpenStudy (anonymous):

no we can do it step by step one third of the runners dropped out 16 of the runners dropped out that leaves 32 now what is three fourths of 32? i.e. what is \(32\times \frac{3}{4}\) ?

OpenStudy (anonymous):

3/8?

OpenStudy (anonymous):

no it should be a whole number 4 goes in to 32 8 times, and 8 times 3 is 24

OpenStudy (anonymous):

i.e. \[32\times \frac{3}{4}=8\times 3=24\] so 24 more runners drop out, leaving \[32-24=8\]

OpenStudy (anonymous):

we could do it another way if you like and get the same answer if one third drop out, then two thirds stay two thirds of 48 is 32 and then if three fourths drop out, one fourth remains one fourth of 32 is 8

OpenStudy (anonymous):

oh cool

OpenStudy (anonymous):

i have another question \[w ^{3}u ^{7}/wu ^{5}\]

OpenStudy (anonymous):

subtract the exponents \[\frac{w^3u^7}{wu^5}=w^{3-1}u^{7-5}=w^2u^2\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

x\[x ^{3}y ^{4}/xy ^{7} \]

OpenStudy (anonymous):

so what would i do if the y for example on the bottom was larger that the top?